Use the method of compensation to show how you can make the subtraction of 563 − 79 easier. 


1
Expert's answer
2021-05-31T15:05:28-0400

Subtraction by compensation means taking away more than you need and then adding some back.

"563 - 79 = (563 + 21) - (79 + 21) = 584 - 100 = 484."